OpenStudy (anonymous):

Hello @jim_thompson5910. I just wanted to get your opinion on if I am correct or not, because these are my corrections. The answers were 2, 3, 7, & 8. But I got 2. The question is below

OpenStudy (anonymous):

If AD = 3x + 5, AB = 7x – 3, and BC = 8x – 10, what is the value of x?

jimthompson5910 (jim_thompson5910):

hopefully you started off with AD = BC ?

OpenStudy (anonymous):

Yes, I did. But I'm not feeling good about my answer at all

jimthompson5910 (jim_thompson5910):

whats your next step?

OpenStudy (anonymous):

3x + 5 = 7x - 3 -4x= -8 x=2 that's what i basically did altogether. the first time i had done this 7x-3=8x-10 -x=-7 x=7 but i did that for AB & BC so of course it was wrong, just not really understanding

jimthompson5910 (jim_thompson5910):

it should be AD = BC

jimthompson5910 (jim_thompson5910):

AD = BC 3x + 5 = 8x – 10 x = ???

OpenStudy (anonymous):

That's 3, right?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

So then I was completely wrong lol. I'm so glad I asked you. Thanks so much. I really appreciate it, you have no idea.

jimthompson5910 (jim_thompson5910):

you're welcome